2. One (1) gram of aradioactive substance disintegrates at the rate of 3.7 × 1010 per


second. The atomic weight of the substance is 226. Calculate its mean life.

1
Expert's answer
2021-12-24T11:32:32-0500

1) Write down the expression for activity of the substance:


"A=\frac NT."


2) Remember what the number of particles is:


"N=nN_a=\frac mMN_a."

3) Express the time:


"T=\frac{N}A=\frac{mN_a}{MA}=\frac{1·6.02·10^{23}}{226·3.7·10^{10}},\\\space\\
T=7.2·10^{10}\text{ s}."